Air entering a dryer has a temperature (dry bulb) of 60°C and a dew point of 26.7°C. Using the humidity chart, determine the actual humidity H, percentage humidity HP, humid heat cS and humid volume vH.
Answer #1

dew point = 26.7*C

% saturation = 100%

From the Humidity chart for 100% humidity

actual humidity at 60 C

H = 0.0225 kg H2O/kg dry air

% Humidity = 14%

Humid heat Cs = 1.005 + 1.88H

= 1.005 + 1.88*0.0225

= 1.047 kJ/kg dry air - K

Humid volume = 2.83*10^-3 T + 4.56*10^-3 T H

= 2.83*10^-3 * (60 +273) + 4.56*10^-3 (60 +273) (0.0225)

= 0.977 m3/kg dry air